When my dad was 31, I was 8. Today my dad is twice as old as I am. How old am I and how old is my dad
Answer by ankor@dixie-net.com(22740)

Let x = Your Dad's age now:
:
"When my dad was 31, I was 8"
:
From this statement we know the age difference of you and your Dad"
31 - 8 = 23 yrs
:
Today my dad is twice as old as I am. How old am I and how old is my dad
x = 2(x-23)
x = 2x - 46
+46 = 2x - x
x = 46 yrs is your Dad now
46 - 23 = 23 yrs is your age
:
:
Check solution in the statement,"my dad is twice as old as I am.
46 = 2(23)